Answer each of the following questions. Assume the asset was purchased at the beginning of the year.

1. A plant asset purchased for $650,000 has an estimated life of 10 years and a residual value of $32,500. Depreciation for the second year of use, determined by the declining-balance method at twice the straight-line rate is $

.
2. A plant asset purchased for $572,000 at the beginning of the year has an estimated life of 5 years and a residual value of $52,000. Depreciation for the third year, determined by the sum-of-the-years'-digits method is $

.
3. A plant asset with a cost of $702,000, estimated life of 5 years, and residual value of $117,000, is depreciated by the straight-line method. This asset is sold for $494,000 at the end of the second year of use. The

lossgain

on the disposal is $

.
Answer #1

1) Depreciation for 2nd year = ($650,000 - 20%) X 20% = $104,000

2) Depreciation for 3rd year = ($572,000 - $52,000) X 3/15 = $104,000

3) Depreciation till 2nd year = ($702,000 - $117,000) X 2/5 = $234,000

Book value at the end of 2nd year = $702,000 - $234,000 = $468,000

Asset is sold for $494,000 which is greater than the book value. Hence there is a gain on sale of asset.

Gain on sale = $494,000 - $468,000 = $26,000